OpenStudy (anonymous):

Consider the sequence 2, 6, 18, 54 .... Let n= the term number in the sequence Let A(n)= The value of the nth term of the sequence 4. What is the common ratio in the sequence? 5. Complete each statement a. A(1) = 2 = 2*3 b. A(2) = 6 = 2*3 = 2*3 c. A(3) = 18 = 2*3*3 = 2*3 d. A(4) = 54 = 2*3*3*3 = 2*3

OpenStudy (raden):

r = a2/a1 = 6/2 = 3 or r = a3/a2 = 18/6 = 3 too

OpenStudy (anonymous):

wait, lol.. what?

OpenStudy (raden):

that's to get the common ratio (r)

OpenStudy (anonymous):

Oh, I forgot to say, I have number four. I need number five. I'm sorry!

OpenStudy (raden):

what do you want for no 5 ? lol, a bit confuse what means " complete statement"

OpenStudy (anonymous):

See, That's my thing. I'm so confused about that. That is what it says in the book. I think they want to know the exponent of each statement.

OpenStudy (raden):

maybe like this : a1 = 2 = 2 * (3)^0 = 2 * 3^(1 - 1) a2 = 6 = 2 * (3)^1 = 2 * 3^(2 - 1) a3 = 18 = 2 * (3)^2 = 2 * 3^(3 - 1) a4 = 54 = 2 * (3)^3 = 2 * 3^(4 - 1) so on ...... we get the general formula of this sequence : an = a1 * r^(n-1)

OpenStudy (raden):

known a1 = 2 and r = 3, so we get an = a1 * r^(n-1) = 2 * 3^(n-1) with a1 = 1st term r = ratio

OpenStudy (anonymous):

I think that's right, You're awesome! Thanks bunches! I'll tell you how I do! :)

OpenStudy (raden):

wlcm :D